OpenStudy (anonymous):

A wheel made 195 revolutions at 14 8/9 revolutions per minute. How many minutes did it rotate.

OpenStudy (shane_b):

\[\frac{195}{14\frac{8}{9}}=?\]

OpenStudy (anonymous):

i have no caculator too solve this problem

OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

i got 1.54 what is the fraction of that

OpenStudy (shane_b):

Convert 14 8/9 into a normal fraction:\[14\frac{8}{9}=\frac{14*9+8}{9}=\frac{134}{9}= about ~ 14.9\]\[\frac{195}{14.9}=13.1minutes\]